40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Fri Jul 24 15:59:12 2020, aaron, Update, Lab Work, cryo open 
    Reply  Fri Jul 24 16:12:13 2020, aaron, Update, Lab Work, cryo open 
Message ID: 2541     Entry time: Fri Jul 24 16:12:13 2020     In reply to: 2540
Author: aaron 
Type: Update 
Category: Lab Work 
Subject: cryo open 

entered W Bridge 264

I stopped by my office (W Bridge 364E) to pick up the new workstation computer and a mini monitor that might be handy. Stayed < 15 min and wore fresh gloves, mask, etc; also recorded entry by door.

I entered the lab around 3:15pm and wiped down the common surfaces. I sanitized and sorted my set of PPE from the box Liz dropped off -- thank you!

We should try not to bring cardboard Into the lab, and remove any cardboard we find. I got rid of a couple boxes.

lab temperature is high

The thermostat is doing its usual thing. It is set to 'cool,' but despite the temperature reading 71.2 F at the thermostat, the temperature at the particle counter is 80F and it feels warm. When I try to lower the set point, the thermostat requests a password. I found online that the default password for is 9995. The thermostat seems to accept the password but won't let me change the setpoint.

The rubber cement at one corner of the HEPA filter delaminated, but the edge held by the kim wipes set in place. I reapplied glue and fixed the tape more securely, and noticed improved air flow through the filter again. The lab temperature also improved a bit -- the leak seems directs cold air down to the floor, while the filter sends cool air across the top of the room.

misc reopening steps

These cover most of the recommended reopening steps. I'll do some general cleaning and organization, and make some new signage, then will call it 'reopen.' 

  • ran the sink hot and cold for about 10 min. There are no noticeable leaks or dampness under the sink, along the edges of the lab, or in any of the major cabinets in the lab.
  • Valved off the cryo Q vacuum and turned off the pump. Pressure reached 2e-3 torr in ~3 hours
  • strain relieved cables of spirou workstation, moved the station, ethernet, etc closer to PSOMA electronics rack
  • checked level of compressed He (1900 psi) and N2 (we don't have a cylinder). I will coordinate with Chub to get an N2 delivery
  • Added isopropyl, methanol, other solvents to the PSOMA inventory
  • Confirmed that our fire extinguisher still has pressure. Where is our eye wash and emergency shower station?  There's one on the basement level, is there one in the subbasement?
  • Not sure about sprinklers, gas, and vacuum utilities. I checked the utility with a yellow valve at NE corner of the lab, but opening the valve releases neither gas nor vacuum. I did not touch the sprinkler system, and did not test the compressed gas tubing (such as for the optics benches).
  • Laser 'in use' lkight at the lab entrance is functioning.
  • Checked that the hot plate (1) and variac cables were not under stress or damaged

exit lab Thu Jul 23 18:09:38 2020

ELOG V3.1.3-